OpenStudy (anonymous):

help!!! 2^n+1 x 2^n-4 / 2^2n-2 :(

OpenStudy (anonymous):

\[2^{n+1}2^{n-4}=2^{n+1+n-4}=2^{2n-3}\] next: \[\frac{2^{2n-3}}{2^{2n-2}}=2^{2n-3-(2n-2)}=2^{-1}\]

OpenStudy (anonymous):

@miszery

OpenStudy (anonymous):

oh okjay so thanksi need to get used to this website, a bit confusing and can't see everything looks so small

OpenStudy (hba):

\[\huge\ \frac{2^{2n-3}}{2^{2n-2}}=2^{2n-3-(2n-2)}=2^{-1}\]

OpenStudy (hba):

^ Now you can see it.

OpenStudy (jiteshmeghwal9):

\[\LARGE{\dfrac{2^{n+1}. 2^{n-4}}{2^{2n-2}}}\]use property \(a^b.a^c=a^{b+c}\) in numerator\[\LARGE{2^{n+1+n-4} \over 2^{2n-2}}\]\[\LARGE{2^{n-3} \over 2^{2n-2}}\]now use the property \(a^b/a^c=a^{b-c}\)\[\LARGE{{2^{n-3-2n+2}=?}}\]

OpenStudy (anonymous):

thanks for the enlargement it's heaps better :)

OpenStudy (anonymous):

ok i got it not thank you... i forgot to change 2n-2 to 2n+2 and got the answer of 2^-5 the first time. cheers :D

OpenStudy (jiteshmeghwal9):

yw, buddy :)

OpenStudy (anonymous):

do i close this now?

OpenStudy (hba):

Yeah,Close this up and post a new question in a new thread if you want :)